SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (July 19, 2005)

Attendees: Rainer Bartoldus, Yaw Ming Chia, Al Eisner, Fang Fang, Selina Li, Brian Lindquist, Dong Su, Dan Walker.

Operations

  • There were a few runs where strange structures were observed in both SVT and DIRC during last Saturday's swing shift. The SVT timestamp distributions showed a slope on top of the normal peak for all sectors while the DIRC timing histograms show a secondary peak around zero with a similar magnitude as the expected peak. This problem wasn't cured by reconfiguring, and it went away after PEP lost the beams. We suspected that this is due to either a beam background problem or an injection problem with possibly missing markers. For more details, see SVT, DIRC, and TRG HN discussions.
  • A dip in TSF SL9 was found since last Sunday's swing shift. See more discussion in DCT/DCZ below.

EMT

  • The .ps file wasn't produced at the end of EMT calibration due to the errors shown in this log file.
  • Dan and Rainer aim to get the EMT prescale in by the end of this month. Rainer asked if there are other system's parameters needed to put in while EMT is loading things into configuration database. This might be a good time to put others in too if there is any. Maybe ZPD input data can also be prescaled?

DCT/DCZ

  • A dip at pivot cell #5 for SL9 showed up in fastmon plots started from run 56246 or so. This problem still persists. See HN posted.
  • Rainer and Su Dong had a meeting with Bill Wisniewski and Walts Inn to talk about getting manpower for solving the TSF problem that involved special TSF power-up sequence. Manchester will draft someone to work on this problem. A possible candidate is Marc. Su Dong suggested that Ming would help cooperate with this as well. This work can be done in the test stand. If a solution is found and well tested, we could implement it during the October down time.

GLT

  • Su Dong got everything working for removing the GLT DAQ delay now, but some more testings are on-going (eg. the board algorithm test). After this has been well tested, the plan is to implement it at IR2 at the end of this month.

L3

  • To avoid L3 node crashes at IR2 (and in SP8), Rainer backed out the changes in the L3 tracking code that was implemented for Run5. See HN posting. This will be in effect in the next online release in for production at IR2. For simulations, this is in effective now.

Data Quality and Monitoring

  • Al suggested us to merge the hbook files to get more statistics for the reference plots.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.